OpenStudy (anonymous):

convert 1024 bits to byte

Parth (parthkohli):

1 byte = 8 bits.

Parth (parthkohli):

Therefore, \(1024 \textbf{ bits} = 1024 \div 8 \textbf{ bytes}\)

Parth (parthkohli):

Easier way to divide it is \(2^{10} \div 2^3\) Now subtract the exponents :D

OpenStudy (anonymous):

1 byte= 8 bits so 1024 bits=1024/8 =128 bytes

OpenStudy (anonymous):

1 byte = 8 bits

OpenStudy (king):

why are u not closing the question?hav u not understood the answer?

OpenStudy (theviper):

\[\Huge{\color{red}{\frac{1024}{8}=128 \checkmark}}\]